OpenStudy (anonymous):

Find the slope of a line that passes through (3, 4) and (-2, 4). Answer undefined 0 eight-fifths five-eighths I think its A. Am I correct?

OpenStudy (anonymous):

yes, it is undefined

OpenStudy (anonymous):

Are you there?

OpenStudy (anonymous):

yes

OpenStudy (ajprincess):

Nope. \(m=\LARGE\frac{y_2-y_1}{x_2-x_1}\) \( =\LARGE\frac{4-4}{-2-3}\) \( =\LARGE\frac{0}{-5}\) \( =0\)

OpenStudy (anonymous):

thank you!

OpenStudy (ajprincess):

yw:)